# Correlated Proxies: A New Definition and Improved Mitigation for Reward Hacking

### v3

@@ C.1 Traffic Control
The traffic control environment, based on the Flow simulator (Wu et al., 2022), simulates a group of human-controlled and RL-controlled vehicles on an on-ramp attempting to merge into traffic on a highway. The true reward prioritizes a small mean commute time, while the proxy reward is the average velocity of all cars. When reward hacking, the RL controlled vehicle on the on-ramp stops indefinitely and lets cars continue forward at high speeds on the highway, which maximizes the proxy reward but increases the commute times of cars on the on-ramp infinitely. As the base policy for the traffic environment we used the Intelligent Driver Model (IDM), a standard approximation of human driving behavior (Treiber et al., 2000). In practice, base policies are often learned via imitation learning, so to simulate this we generate data from the IDM controller and train a behavioral cloning (BC) policy using the generated data.
@@ C.2 Glucose Monitoring
The RL agent (bottom) controls the insulin administered to a simulated patient in order to maintain healthy glucose levels. The true reward is a standard measure of health risk for the patient, but the proxy reward prioritizes the monetary cost of insulin. As the safe baseline policy, we train a BC policy based on data generated by a PID controller with parameters tuned by the original designers of the simulator (Steil, 2013).
@@ C.3 Pandemic Mitigation
The RL agent chooses the level of lockdown restrictions placed on the population by observing the results of testing. The proxy reward function omits the political cost associated with certain decisions. Our base policy is trained via BC on a combination of hand-specified and real-world strategies employed by governments during the pandemic, which were also used by Kompella et al. (2020) as baselines.
@@ C.4 RLHF
We base our RLHF environment on the work of Coste et al. (2024), who study overoptimization of LLM-based reward models. The proxy reward model we use is fine-tuned from Pythia-70M (Biderman et al., 2023) on the AlpacaFarm (Dubois et al., 2023) preference dataset. As a true reward model, we use the Llama 3 Tulu V2 8B RM from AI2 (Ivison et al., 2024). As a base policy, we use Coste et al.’s SFT policy, which was fine-tuned from Pythia-1.4B on the AlpacaFarm SFT data. We evaluate policies’ true and proxy rewards on responses sampled for a held out set of prompts.
@@ C.5 Tomato-Watering Gridworld
The tomato environment contains a sprinkler state where the agent perceives all tomatoes as being watered and thus receives high proxy reward but no true reward. We train a base policy using the true reward function, and then add a 10% chance of taking a random action to ensure there is room to improve upon it.
@@ D.1 Non-LLM Experiments
Initializing using an imitation learning policy has been shown to effectively speed up the learning process (Laidlaw et al., 2023; Uchendu et al., 2023) and is used in practice for RLHF (Stiennon et al., 2020), so we initialize our policies using the specified $\pi_{\text{base}}$} for the more realistic traffic, glucose, and pandemic environments.
@@ D.2 RLHF Experiments
$\chi^{2}$ divergence: $\lambda\left(\frac{\pi_{\theta}(a\mid s)}{{\pi_{\text{base}}}(a\mid s)}+\frac{{\pi_{\text{base}}}(a\mid s)}{\pi_{\theta}(a\mid s)}-2\right)$
KL divergence: $\lambda\left(\log\frac{\pi_{\theta}(a\mid s)}{{\pi_{\text{base}}}(a\mid s)}+\frac{{\pi_{\text{base}}}(a\mid s)}{\pi_{\theta}(a\mid s)}-1\right)$
where $s$ is the prompt and $a$ is the sampled response. Intuitively, both loss terms have a unique minimum when $\pi_{\theta}(a\mid s)={\pi_{\text{base}}}(a\mid s)$ and in expectation are equivalent to the correct divergence. Schulman (2020) suggests that these are particularly low-variance estimates, and we find that they work well in practice.

### v4

@@ C.1 Traffic Control
The traffic control environment, based on the Flow simulator (Wu et al., 2022), simulates a group of human-controlled and RL-controlled vehicles on an on-ramp attempting to merge into traffic on a highway. The true reward prioritizes a small mean commute time, while the proxy reward is the average velocity of all cars. When reward hacking, the RL controlled vehicle on the on-ramp stops indefinitely and lets cars continue forward at high speeds on the highway, which maximizes the proxy reward but increases the commute times of cars on the on-ramp infinitely. As the reference policy for the traffic environment we used the Intelligent Driver Model (IDM), a standard approximation of human driving behavior (Treiber et al., 2000). In practice, reference policies are often learned via imitation learning, so to simulate this we generate data from the IDM controller and train a behavioral cloning (BC) policy using the generated data.
@@ C.2 Glucose Monitoring
The RL agent (bottom) controls the insulin administered to a simulated patient in order to maintain healthy glucose levels. The true reward is a standard measure of health risk for the patient, but the proxy reward prioritizes the monetary cost of insulin. As the safe baseline policy, we train a BC policy using data generated by a PID controller with parameters tuned by the original designers of the simulator (Steil, 2013).
@@ C.3 Pandemic Mitigation
The RL agent chooses the level of lockdown restrictions placed on the population by observing the results of testing. The proxy reward function omits the political cost associated with certain decisions. Our reference policy is trained via BC on a combination of hand-specified and real-world strategies employed by governments during the pandemic, which were also used by Kompella et al. (2020) as baselines.
@@ C.4 RLHF
We base our RLHF environment on the work of Coste et al. (2024), who study overoptimization of LLM-based reward models. The proxy reward model we use is fine-tuned from Pythia-70M (Biderman et al., 2023) on the AlpacaFarm (Dubois et al., 2023) preference dataset. As a true reward model, we use the Llama 3 Tulu V2 8B RM from AI2 (Ivison et al., 2024). As a reference policy, we use Coste et al.’s SFT policy, which was fine-tuned from Pythia-1.4B on the AlpacaFarm SFT data. We evaluate policies’ true and proxy rewards on responses sampled for a held out set of prompts.
@@ C.5 Tomato-Watering Gridworld
The tomato environment contains a sprinkler state where the agent perceives all tomatoes as being watered and thus receives high proxy reward but no true reward. We train a reference policy using the true reward function, and then add a 10% chance of taking a random action to ensure there is room to improve upon it.
@@ D.1 Non-LLM Experiments
Initializing using an imitation learning policy has been shown to effectively speed up the learning process (Laidlaw et al., 2023; Uchendu et al., 2023) and is used in practice for RLHF (Stiennon et al., 2020), so we initialize our policies using the specified $\pi_{\text{ref}}$} for the more realistic traffic, glucose, and pandemic environments.
@@ D.2 RLHF Experiments
$\chi^{2}$ divergence: $\lambda\left(\frac{\pi_{\theta}(a\mid s)}{{\pi_{\text{ref}}}(a\mid s)}+\frac{{\pi_{\text{ref}}}(a\mid s)}{\pi_{\theta}(a\mid s)}-2\right)$
KL divergence: $\lambda\left(\log\frac{\pi_{\theta}(a\mid s)}{{\pi_{\text{ref}}}(a\mid s)}+\frac{{\pi_{\text{ref}}}(a\mid s)}{\pi_{\theta}(a\mid s)}-1\right)$
where $s$ is the prompt and $a$ is the sampled response. Intuitively, both loss terms have a unique minimum when $\pi_{\theta}(a\mid s)={\pi_{\text{ref}}}(a\mid s)$ and in expectation are equivalent to the correct divergence. Schulman (2020) suggests that these are particularly low-variance estimates, and we find that they work well in practice.